# Evaluating LLM-driven User-Intent Formalization for Verification-Aware Languages

### v1

@@ Abstract
Recent work has developed a framework for evaluating and benchmarking the user-intent formalization problem for mainstream programming languages [12]. However, as we argue in this paper, such an approach does not readily extend to verification-aware languages that support rich specifications (using quantifiers and ghost variables) that cannot be evaluated through dynamic execution. Previous work also required generating program mutants using LLMs to create the benchmark. We advocate an alternate, perhaps simpler approach of symbolically testing specifications to provide an intuitive metric for evaluating the quality of specifications that can be easily instantiated with most verification-aware languages. We demonstrate that our automated metric agrees closely with a GPT-4 generated and human-labeled dataset of roughly 150 Dafny specifications for the popular MBPP code-generation benchmark, yet demonstrates cases where the human labeling is not perfect. We believe our work provides a stepping stone to enable the establishment of a benchmark and research agenda for the problem of user-intent formalization for programs.
@@ I Introduction
Large language models (LLMs) have recently demonstrated potential to bridge the gap between informal intent and formal artefact such as code [9, 5], by performing code generation from natural language. Such progress has been spurred in large parts through the creation of crowd-sourced benchmarks such as HumanEval [9] and Mostly Basic Python Programs (MBPP) [5] with automated program-semantics based metrics (such as tests), unlike NLP metrics such as BLEU scores [24]. These benchmarks allow a community to evaluate their models and techniques on a common set of benchmarks automatically and measure progress. One can hope that the establishment of similar benchmarks and metrics for specification generation (decoupled from code generation) can enable synthesizing useful specifications from informal intent in practice.
@@
Motivated by such a need, Endres et al. [12] describe the problem of generating formal declarative specifications (namely, method postconditions) from informal intent using LLMs and automated metrics for evaluating them — we refer to this as user-intent formalization problem in this paper. For code generation, the quality of an implementation generated from informal intent is measured through the set of hidden validation tests (that the code generator does not have access to). For user-intent formalization, Endres et al. re-purpose code-generation benchmarks (such as HumanEval) and introduce two metrics (a) correctness and (b) completeness, with respect to the set of (hidden) validation tests and the (hidden) reference implementation. Correctness captures that the specification satisfies the reference implementation for all the tests, and completeness captures the power of the specification to discriminate against buggy mutations of the reference implementation (given the set of tests). Further, these mutants were generated through LLMs and are grouped by the subset of tests that fail them. They demonstrate that these automatic metrics closely resemble the quality of specifications as determined through manual analysis. In this paper, we investigate creating benchmarks (and associated metrics) for user-intent formalization for verification-aware languages such as F* [26], Dafny [21] and Verus [20] etc. A verification-aware language supports a rich program logic for expressing specifications and offers the ability to verify them statically using automated theorem provers. However, for most non-trivial programs, the verification requires manually decomposing the problem through the use of ghost variables, intermediate lemmas, invariants and assertions. We focus on the problem of synthesizing and evaluating specifications automatically.
I-A Running example
@@
Although the implementation satisfies the specification, there is no evidence that the specification indeed captures what the user intended (and specified implicitly in the hidden test cases). The authors perform manual review of the specifications and mark them as either to denote if the specification is respectively, inconsistent, weakly or strongly consistent with the intent expressed in the tests [2]. Therefore, although this work addresses generating verified Dafny programs from informal intent, it relies completely on a human to ensure that the specification matches the intent. In particular, the framework allows for the Dafny code to only satisfy a vacuous specification (such as ensures true) or worse, an incorrect one. This aspect renders this dataset not suited for a benchmark for code generation in verification-aware languages, as any alternate specification will need to be labeled by a human, which makes it subjective as well as non-automated.
I-B Proposal: Symbolically testing specifications
@@
We first argue that the none of the prior approaches are readily applicable to yield an automatic metric: First, running the tests on the implementation without specifications will not rule out vacuous specifications. Second, one cannot apply the approach of Endres et al. for rich specifications that contain ghost state or constructs that cannot be evaluated at runtime (such as a universal/existential quantifier). One can ignore tests and attempt to statically verify the synthesized specification directly against the hidden reference implementation. This can still allow vacuous specifications such as true. Further, it is unlikely that such verification would be automatic for non-trivial specifications given the need to infer intermediate lemmas and invariants. One can instead provide a hidden reference specification and check the candidate specification for semantic equivalence. However, such a metric would be too strict as it would not be able to distinguish weak and vacuous specifications (such as true) from strong (yet incomplete) specifications.
Next, we propose a method for evaluating user-intent formalization for verification-aware languages that leverages the validation tests and symbolic verification capabilities of these languages. Given a set of tests denoting the a set of consistent input-output pairs, we define the correctness and completeness metrics purely over the tests, without the need for an implementation.
@@ I-B1 Correctness
I-B2 Completeness
The completeness measure for a specification given T is the fraction of output mutations of the tests in T that is inconsistent with. Let be a finite set of mutations of T. In this work, we restrict the set of mutants per input to a fixed number (5 for this paper). Let be the largest subset such that for each , the following Hoare-triple does not hold:
Then the completeness measure of with respect to T is (this is inspired by kill-set in mutation-testing literature [16]). For curious readers, we also contrast these with a plausible proposal in Appendix B.
Let us demonstrate an implementation of these Hoare triples as Dafny programs for our running example. For correctness against the first test (in JSON), we create the following Dafny program by providing the body of SharedElements:
The Dafny program symbolically checks (using SMT solvers [6]) that the specification holds for the specific input and output. The two assume statements constrain the content of input parameters a and b to equal values in a1 and a2 respectively. The redundant asserts are needed for the verifier to trigger the quantifiers used in InArray to enable the proof. Finally, the output variable is assigned one of the expected values ([5,4] is also an acceptable value).
The above verification only proves that the specification is correct for the test; a vacuous specification ensures true would also be verified. For completeness, we mutate the output value in result in several ways and check that the verification fails. In this case, mutating the value in result to [6] would fail the first postcondition, since 6 is not present in both the input arrays.
@@ I-B2 Completeness
One may caution against using the verification failure (as used for the completeness checks) as a means to show that the corresponding Hoare-triple does not hold. In other words, given that deductive verifiers are sound, but not necessarily complete, a verification failure is typically interpreted as an unknown outcome. However, note that we check for completeness for a specification against an input-output example only if is proved correct for all the tests in T (including the original test ). Since the two verification conditions differ only in a concrete value (between and ), we conjecture that the underlying reasoning (including quantifiers instantiated) would be quite identical in most cases, and the verification failure strongly indicates that the Hoare-triple does not hold.
@@ II Implementation and evaluation
We report ongoing work in implementing the two metrics for the MBPP-DFY dataset mentioned earlier in Section I-A. Our tool (a 400 line Python script) consumes the method signature, test cases and the candidate specification from the JSON and Dafny files, and creates Dafny programs for verifying the correctness and completeness for each test. It then invokes the Dafny verifier, and reports the aggregate correctness as well completeness score for each specification, averaged over the different test cases; it finally compares the metrics against the labels provided by authors. Of the 153 problems with specifications (a subset written by humans), we have currently managed to apply our tool to check 64 of the specifications.
@@ II-B Results
We find that for large majority of examples, the manual labels align with the metrics computed by our tool. That is, for the incorrect specifications, our tool reports a verification failure for correctness check, and report a high completeness score (usually ) for precise specifications. This provides evidence that the automated metrics serve as a good proxy for the user label. One such incorrect specification generated by GPT-4 (and correctly labeled) is the problem of RemoveDuplicates (”task_id” 572) (also the motivating example by Endres et al. [12] demonstrating ambiguity of natural language) where the task is to remove all elements with duplicates, but GPT-4 interprets it as the problem of retaining a single copy of each value. Similarly, for countSubstrings (”task_id” 61), that counts the number of substrings whose sum equals their length, our completeness check gives a low score to the specification (correctly labeled weak_spec) that only ensures count is a non-negative number.
@@
Interestingly, we also discovered a few cases where the correctness check failed for specifications labeled strong_spec. These include examples with ”task_id”s 234, 240, and 445. For these examples, the authors had accidentally introduced bugs while while copying from the test cases in Python. For instance, for ”task_id” 234 that computes the cube of a number, the authors mistakenly copied the value 25 instead of 125 for the cube of 5. We noticed that a couple of these errors have been fixed in the latest commit at the time of writing (fb4f53e), but a few still remain (e.g., 445 where the value 32 is replaced by 31).
@@ II-C Limitations
We have already automated a few failed proofs due to quantifier instantiation by adding intermediate redundant assertions equating two arrays on every index (see assert statements in the body of SharedElements in Sec I-A). We are currently automating specifying the amount of fuel and quantifier instantiation witnesses when ranging over the indices of an array.
III Related work and conclusion
Although the literature on specification mining is fairly rich, it has historically focused on techniques that infer specifications of an implementation through dynamic [13] and recently neural techniques [25]. The objective of these techniques is to learn invariants from a few runs that generalize to unseen executions. Static [14, 19, 23] and more recently neural approaches [27, 17] have also been applied to the problem (inductive) invariant generation to aid program proofs. On the other hand, work on user-intent formalization (i.e., translating natural language comments to specifications) [7, 12] synthesize the intended specification, even in the absence of an implementation. The problem of generating test oracle assertions have also been investigated [11], but they only apply to single input or test prefix. TiCoder [18] introduced the term ”user-intent formalization” over test cases for interactive code generation, and alluded to metrics for measuring quality of weak specifications (namely, tests) in terms of correctness (user acceptance) and completeness (prunes buggy codes). The work on autoformalization [28] for translating natural language comments to mathematical theorems is closest to our work. However, these techniques do not use semantic checks (such as tests and symbolic verification) to ensure the quality of the generated formal mathematical statements; instead they leverage textual similarity metrics such as BLEU.
Based on the encouraging results so far, we are optimistic to create a benchmark with the metrics with a large fraction of examples from MBPP-DFY dataset. Just as the creation of MBPP and HumanEval has turbo-charged progress on high-quality code generation, we hope that the benchmark will accelerate research and practice of synthesizing high-quality specifications from informal intent in verification-aware languages. This when coupled with work on program and proof synthesis (given a specification) can greatly lower the cost required to create formally verified modules in the future [8].
@@ Appendix B Alternate proposal
A reader may wonder if the following check achieves a similar objective:

### v2

@@ Abstract
Recent work has developed a framework for evaluating and benchmarking the user-intent formalization problem for mainstream programming languages [12]. However, as we argue in this paper, such an approach does not readily extend to verification-aware languages that support rich specifications (using quantifiers and ghost variables) that cannot be evaluated through dynamic execution. Previous work also required generating program mutants using LLMs to create the benchmark. We advocate an alternate, perhaps simpler approach of symbolically testing specifications to provide an intuitive metric for evaluating the quality of specifications that can be easily instantiated with most verification-aware languages. We demonstrate that our automated metric agrees closely on a human-labeled dataset of Dafny specifications for the popular MBPP code-generation benchmark, yet demonstrates cases where the human labeling is not perfect. We also outline formal verification challenges that need to be addressed to apply the technique more widely. We believe our work provides a stepping stone to enable the establishment of a benchmark and research agenda for the problem of user-intent formalization for programs.
@@ I Introduction
Large language models (LLMs) have recently demonstrated potential to bridge the gap between informal intent and formal artefact such as code [9, 5], by performing code generation from natural language. Such progress has been spurred in large parts through the creation of crowd-sourced benchmarks such as HumanEval [9] and Mostly Basic Python Programs (MBPP) [5] with automated program-semantics based metrics (such as tests), unlike NLP metrics such as BLEU scores [25]. For code generation, the quality of an implementation generated from informal intent is measured through the set of hidden validation tests. We term the tests as ”hidden” since they are not available to the model at the time of code generation. The benchmarks also provide a reference code in Python for each problem, which are also hidden from the code generation model. These approaches rely heavily on the presence of a high-quality set of validation tests to exercise most corner cases. These benchmarks allow a community to evaluate their models and techniques on a common set of benchmarks automatically and measure progress. One can hope that the establishment of similar benchmarks and metrics for specification generation (decoupled from code generation) can enable synthesizing useful specifications from informal intent in practice.
@@
Motivated by such a need, Endres et al. [12] describe the problem of generating formal declarative specifications (namely, method postconditions) from informal intent using LLMs and automated metrics for evaluating them — we refer to this as user-intent formalization problem in this paper. For user-intent formalization, Endres et al. re-purpose code-generation benchmarks (such as HumanEval) and introduce two metrics (a) correctness and (b) completeness, with respect to the set of (hidden) validation tests and the (hidden) reference code. Correctness captures that the specification satisfies the reference code for all the validation tests; completeness captures the strength of the specification to discriminate against buggy mutations of the reference code (under the set of tests). Further, these code mutants are generated through LLMs and are grouped by the subset of tests that fail them. They demonstrate that these automatic metrics closely resemble the quality of specifications as determined through manual analysis. As with code generation, the approach relies on the presence of a high-quality test-suite.
I-A Summary
In this paper, we investigate creating benchmarks (dataset and associated automatic metrics) for user-intent formalization for verification-aware languages such as F* [27], Dafny [21] and Verus [20]. A verification-aware language supports a rich program logic for expressing specifications and offers the ability to verify them statically using automated theorem provers. However, for most non-trivial programs, the verification requires manually decomposing the problem through the use of ghost variables, intermediate lemmas, invariants and assertions.
We focus on the problem of evaluating specifications automatically. We discuss why prior approaches do not readily apply to evaluating specifications. Then we propose the use of automatic program verification to symbolically ”test” the specifications to determine their quality. We have developed a prototype and applied it to a dataset of Dafny specifications from prior work [23]. We demonstrate that the quality of the specifications obtained through the automated means aligns well with the human-labeling of the specifications in most cases. Finally, we describe the unique challenges (such as quantifier instantiation) that need to be addressed to apply the technique to a larger class of problems.
I-B Running example
@@
Although the implementation satisfies the specification, there is no evidence that the specification indeed captures what the user intended (and specified implicitly in the hidden test cases). The authors perform manual review of the specifications and mark them as either to denote if the specification is respectively, inconsistent, weakly or strongly consistent with the intent expressed in the natural language and tests [2]. Therefore, although this work addresses generating verified Dafny programs from informal intent, it relies completely on a human to ensure that the specification matches the intent. In particular, the framework allows for the Dafny code to only satisfy a vacuous specification (such as ensures true) or worse, an incorrect one. This aspect renders this dataset unsuitable as an automated benchmark for specification generation in verification-aware languages. In other words, if a language model generates a specification that is non-equivalent to the ground-truth specification labeled by the user, we cannot determine if it is incorrect without requiring a user. This also makes the evaluation subjective. Since verification of (generated) Dafny code is only as good as the specification, this makes the dataset unsuitable for an automated benchmark for verified code generation as well.
I-C Proposal: Symbolically testing specifications
@@
We first argue that the none of the prior approaches are readily applicable to yield an automatic metric: First, running the tests on the implementation without specifications will not rule out vacuous specifications. Second, one cannot apply the approach of Endres et al. for rich specifications that contain ghost state or constructs that cannot be evaluated at runtime (such as a universal/existential quantifier). One can ignore tests and attempt to statically verify the synthesized specification directly against the hidden reference code. This can still allow vacuous specifications such as true. Further, it is unlikely that such verification would be automatic for non-trivial specifications given the need to infer intermediate lemmas and invariants. One can instead provide a hidden reference specification and check the candidate specification for semantic equivalence. However, such a metric would be too strict as it would not be able to distinguish weak and vacuous specifications (such as true) from strong (yet incomplete) specifications.
Next, we propose a method for evaluating user-intent formalization for verification-aware languages that leverages the validation tests and symbolic verification capabilities of these languages. Given a set of tests consisting of a set of consistent input-output pairs, we define the correctness and completeness metrics purely over the tests, without the need for the reference code. Of course, just as in the case of code generation, the technique for determining specification quality assumes a high-quality set of validation tests.
@@ I-C1 Correctness
I-C2 Completeness
The completeness measure for a specification given T is the fraction of output mutations of the tests in T that is inconsistent with. Let be a finite set of mutations of T that mutate the output values for the given inputs. In this work, we restrict the set of mutants per input to a fixed number (5 for this paper). Let be the largest subset such that for each , the following Hoare-triple does not hold:
Then the completeness measure of with respect to T is (this is inspired by kill-set in mutation-testing literature [16]). For interested readers, we also contrast these with a plausible proposal in Appendix B.
Let us demonstrate an implementation of these Hoare triples as Dafny programs for our running example. Consider the first test (see JSON input in Section I-B) that asserts that SharedElements should return the set for input arrays and . For correctness against this test, we create the following Dafny program by providing a definition of SharedElements that performs the soundness check:
The Dafny program above is identical to the program in Section I-B, except for the body of SharedElements specified within the curly braces . Instead of the original implementation of the method, we model the Hoare-triple for correctness described above. The precondition of the Hoare-triple (true) translates to assume true which is dropped. The input assignment is modeled as assignments of the input to temporary variables a1, a2, followed by constraining the actual parameters a, b respectively. For Dafny, this amounts to saying that two arrays a (respectively, b) and a1 (respectively, a2) are equal on all elements up to their lengths (which implies that the lengths are identical as well). The redundant asserts are needed for the verifier to trigger the quantifiers used in InArray to enable the proof. Finally, the output parameter result is assigned one of the expected values ([5,4] is also an acceptable value). The Dafny program symbolically checks (using Satisfiability Modulo Theories solvers [6]) that the specification (provided by the ensures statements) holds for the specific input and output.
However, the above verification only proves that the specification is correct for the test; a vacuous specification ensures true would also be verified. For completeness, we mutate the output value in result in several ways and check that the verification fails. In this case, mutating the value in result to [6] would fail the first postcondition, since 6 is not present in both the input arrays.
@@ I-C2 Completeness
One may caution against using a verification failure (as used for the completeness checks) as a means to show that the corresponding Hoare-triple does not hold. In other words, given that deductive verifiers are sound, but not necessarily complete, a verification failure is typically interpreted as an unknown outcome. However, note that we check for completeness for a specification against an input-output example only if is proved correct for all the tests in T (including the original test ). Since the two verification conditions differ only in a concrete value (between and ), we conjecture that the underlying reasoning (including quantifiers instantiated) would be quite identical in most cases, and the verification failure strongly indicates that the Hoare-triple does not hold.
@@ II Implementation and evaluation
We report ongoing work in implementing the two metrics for the MBPP-DFY dataset mentioned earlier in Section I-B. Our tool (a 400 line Python script) consumes the method signature, test cases and the candidate specification from the JSON and Dafny files, and creates Dafny programs for verifying the correctness and completeness for each test. It then invokes the Dafny verifier, and reports the aggregate correctness as well completeness score for each specification, averaged over the different test cases; it finally compares the metrics against the labels provided by authors. Of the 153 problems with specifications (a subset written by humans), we have managed to apply our tool to check 64 of the specifications (at the time of writing). In other words, for these 64 examples, our tool is able to verify the soundness of the specifications over the set of validation tests. We have released the scripts, dataset and outputs at the website https://github.com/microsoft/nl-2-postcond.
@@ II-B Results
We find that for large majority of examples, the manual labels align with the metrics computed by our tool. That is, for the incorrect specifications, our tool reports a verification failure for correctness check, and report a high completeness score (usually ) for precise specifications. In other words, barring a few exceptions below, all the specifications that are marked strong_spec have a completeness score above , and all the weak_spec have a score below .
This provides evidence that the automated metrics serve as a good proxy for the user label. One such incorrect specification generated by GPT-4 (and correctly labeled) is the problem of RemoveDuplicates (”task_id” 572) (also the motivating example by Endres et al. [12] demonstrating ambiguity of natural language) where the task is to remove all elements with duplicates, but GPT-4 interprets it as the problem of retaining a single copy of each value. Similarly, for countSubstrings (”task_id” 61), that counts the number of substrings whose sum equals their length, our completeness check gives a low score to the specification (correctly labeled weak_spec) that only ensures count is a non-negative number.
@@
Interestingly, we also discovered a few cases where the correctness check failed for specifications labeled strong_spec. These include examples with ”task_id”s 234, 240, and 445. For these examples, the authors had accidentally introduced bugs while copying from the test cases in Python. For instance, for ”task_id” 234 that computes the cube of a number, the authors mistakenly copied the value 25 instead of 125 for the cube of 5. We noticed that a couple of these errors have been fixed in the latest commit at the time of writing (fb4f53e), but a few still remain (e.g., 445 where the value 32 is replaced by 31).
@@ II-C Limitations
We have already automated a few failed proofs due to quantifier instantiation by adding intermediate redundant assertions equating two arrays on every index (see assert statements in the body of SharedElements in Sec I-B). In future work, we plan to automate the inference of fuel amount and quantifier instantiation witnesses when ranging over the indices of an array.
III Related work
Although the literature on specification mining is fairly rich, it has historically focused on techniques that infer specifications of an implementation through dynamic [13] and recently neural techniques [26]. The objective of these techniques is to learn invariants from a few runs that generalize to unseen executions. Static [14, 19, 24] and more recently neural approaches [28, 17] have also been applied to the problem (inductive) invariant generation to aid program proofs. Finally, recent works have explored the generation of programs, specifications and proofs from natural languages [23, 29, 22]. On the other hand, work on user-intent formalization (i.e., translating natural language comments to specifications) [7, 12] synthesize the intended specification, even in the absence of an implementation. The problem of generating test oracle assertions have also been investigated [11], but they only apply to single input or test prefix. TiCoder [18] introduced the term ”user-intent formalization” over test cases for interactive code generation, and alluded to metrics for measuring quality of weak specifications (namely, tests) in terms of correctness (user acceptance) and completeness (prunes buggy codes). The work on autoformalization [30] for translating natural language comments to mathematical theorems is closest to our work. However, these techniques do not use semantic checks (such as tests and symbolic verification) to ensure the quality of the generated formal mathematical statements; instead they leverage textual similarity metrics such as BLEU.
IV Conclusion
In this paper, we motivate the problem of evaluating user-intent-formalization for verification-aware languages. We demonstrate that the idea of symbolically testing specifications against validation tests can provide an automated metric for a benchmark. We plan to curate a benchmark with a large fraction of examples from MBPP-DFY dataset using the above metric. We hope such benchmarks will accelerate the research on specification generation from informal intent in verification-aware languages. This when coupled with work on program and proof synthesis (given a specification) can greatly lower the cost required to create formally verified modules in the future [8].
@@ Appendix B Alternate proposal
A reader may wonder if the following check achieves a similar objective for checking completeness: